# AMSP-UOD: When Vortex Convolution and Stochastic Perturbation Meet Underwater Object Detection

### 14 Dec 2023 v2

@@ Abstract
In this paper, we present a novel Amplitude-Modulated Stochastic Perturbation and Vortex Convolutional Network, AMSP-UOD, designed for underwater object detection. AMSP-UOD specifically addresses the impact of non-ideal imaging factors on detection accuracy in complex underwater environments. To mitigate the influence of noise on object detection performance, we propose AMSP Vortex Convolution (AMSP-VConv) to disrupt the noise distribution, enhance feature extraction capabilities, effectively reduce parameters, and improve network robustness. We design the Feature Association Decoupling Cross Stage Partial (FAD-CSP) module, which strengthens the association of long and short range features, improving the network performance in complex underwater environments. Additionally, our sophisticated post-processing method, based on Non-Maximum Suppression (NMS) with aspect-ratio similarity thresholds, optimizes detection in dense scenes, such as waterweed and schools of fish, improving object detection accuracy. Extensive experiments on the URPC and RUOD datasets demonstrate that our method outperforms existing state-of-the-art methods in terms of accuracy and noise immunity. AMSP-UOD proposes an innovative solution with the potential for real-world applications. Our code is available at https://github.com/zhoujingchun03/AMSP-UOD.
@@ Introduction
In this paper, we propose the AMSP-UOD network, crafted to tackle non-ideal imaging factors in underwater environments. Utilizing the optical imaging model $I=H(J,B,t)+N$ ($I$ represent observed image, $J$ represent raw scene, $B$ represents backscatter, and $t$ represent transmission map), we discern that underwater images combine a degradation function $H$ with noise $N$. To remove noise, we propose the AMSP-VConv. This strategy not only reduces parameters but also bolsters the network’s robustness. We further implement the FAD-CSP to improve feature extraction in degraded environments. Our post-processing strategy, which relies on NMS, is designed to optimize the detection of dense clusters of underwater objects. Experimental results on the URPC and RUOD datasets showcase the effectiveness of our method. Overall, AMSP-UOD presents an innovative solution for UOD with potential real-world applications.
@@
Global Feature-Aware Representation
In convolutional operations for global feature processing, a deeper network structure is usually required to extract rich feature information. This often increases the likelihood of the network getting trapped in local optima. To address this issue, we devised an efficient global feature-aware module and seamlessly integrated it into the FAD-CSP network using an attention mechanism. The structure of the proposed global feature-aware module is depicted in Figure 3. For a given input tensor $F_{in}$, we processing it through a bar-shaped pooling group (Equations (16) and (17)), which compresses salient features into a one dimensional space. This method not only trades longer distance feature correlations, but also exhibits much lower computational overhead compared to convolution. This process can be expressed as follows:
@@
Figure 5: Visualization of object detection results of different object detection methods on URPC (ZhanJiang). (a) YOLOv3 (Redmon and Farhadi 2018), (b) YOLOv5s (Jocher et al. 2020), (c) YOLOv6s(Li et al. 2022), (d) YOLOv7-tiny (Wang, Bochkovskiy, and Liao 2023), (e) Faster R-CNN (Girshick 2015), (f) Cascade R-CNN (Cai and Vasconcelos 2018), (g) RetinaNet (Lin et al. 2017), (h) FCOS (Tian et al. 2019), (i) ATSS (Zhang et al. 2020b), (j) TOOD (Feng et al. 2021), (k) PAA (Kim and Lee 2020), (l) Ours-Standard, (m) Ours-AMSP-VConv, (n) Ours-AMSP-VConv + NMS Similar.
@@
We adopt AP and AP50 as the primary metrics for model accuracy evaluation, with precision (P) and recall (R) as supplementary indicators. To showcase the generalizability of our network, we trained it on the URPC (ZhanJiang) dataset, from the 2020 National Underwater Robotics Professional Competition and the extensive RUOD dataset. The URPC dataset contains 5,543 training images across five categories, with 1,200 images from its B-list answers serving as the test set. RUOD, with various underwater scenarios, consists of 10 categories, with 9,800 training images and 4,200 test images.
@@
Table 3: Comparison with existing methods on the URPC and RUOD datasets. Ho: holothurian’s AP50, Ec: echinus’s AP50, St: starfish’s AP50, Sc: scallop’s AP50, Wa: waterweeds’s AP50. AP: AP@[0.5:0.05:0.95], AP50: AP@0.5. Bold is highest, underlined is second-highest.
@@
Ablation of AMSP-VConv
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (Nos. 62301105 and 61702074), the 2022 National Undergraduate Innovation and Entrepreneurship Training Program Project (No.202210577003), National Key Research and Development Program of China (No.2018AAA0100400), China Postdoctoral Science Foundation (NO.2021M701780), the High Performance Computing Center of Dalian Maritime University, and the Supercomputing Center of Nankai University. We are also sponsored by CAAI-Huawei MindSpore Open Fund.

### 18 Jan 2024 v3

@@ Abstract
In this paper, we present a novel Amplitude-Modulated Stochastic Perturbation and Vortex Convolutional Network, AMSP-UOD, designed for underwater object detection. AMSP-UOD specifically addresses the impact of non-ideal imaging factors on detection accuracy in complex underwater environments. To mitigate the influence of noise on object detection performance, we propose AMSP Vortex Convolution (AMSP-VConv) to disrupt the noise distribution, enhance feature extraction capabilities, effectively reduce parameters, and improve network robustness. We design the Feature Association Decoupling Cross Stage Partial (FAD-CSP) module, which strengthens the association of long and short range features, improving the network performance in complex underwater environments. Additionally, our sophisticated post-processing method, based on Non-Maximum Suppression (NMS) with aspect-ratio similarity thresholds, optimizes detection in dense scenes, such as waterweed and schools of fish, improving object detection accuracy. Extensive experiments on the URPC and RUOD datasets demonstrate that our method outperforms existing state-of-the-art methods in terms of accuracy and noise immunity. AMSP-UOD proposes an innovative solution with the potential for real-world applications. Our code is available at: https://github.com/zhoujingchun03/AMSP-UOD.
@@ Introduction
In this paper, we propose the AMSP-UOD network, crafted to tackle non-ideal imaging factors in underwater environments. Utilizing the optical imaging model $I=H(J,B,t)+N$ ($I$ represent observed image, $J$ represent raw scene, $B$ represents backscatter, and $t$ represent transmission map), we discern that underwater images combine a degradation function $H$ with noise $N$. To remove noise, we propose the AMSP-VConv. This strategy not only reduces parameters but also bolsters the network’s robustness. We further implement the FAD-CSP to improve feature extraction in degraded environments. Our post-processing strategy, which relies on NMS, is designed to optimize the detection of dense clusters of underwater objects. Experimental results on the URPC (Liu et al. 2021) and RUOD (Fu et al. 2023) datasets showcase the effectiveness of our method. Overall, AMSP-UOD presents an innovative solution for UOD with potential real-world applications.
@@
Global Feature-Aware Representation:
In convolutional operations for global feature processing, a deeper network structure is usually required to extract rich feature information. This often increases the likelihood of the network getting trapped in local optima. To address this issue, we devised an efficient global feature-aware module and seamlessly integrated it into the FAD-CSP network using an attention mechanism. The structure of the proposed global feature-aware module is depicted in Figure 3. For a given input tensor $F_{in}$, we processing it through a bar-shaped pooling group, which compresses salient features into a one dimensional space. This method not only trades longer distance feature correlations, but also exhibits much lower computational overhead compared to convolution. This process can be expressed as follows:
@@
Figure 5: Visualization of object detection results of different object detection methods on URPC (Zhanjiang). (a) YOLOv3 (Redmon and Farhadi 2018), (b) YOLOv5s (Jocher 2020), (c) YOLOv6s(Li et al. 2022), (d) YOLOv7-tiny (Wang, Bochkovskiy, and Liao 2023), (e) Faster R-CNN (Girshick 2015), (f) Cascade R-CNN (Cai and Vasconcelos 2018), (g) RetinaNet (Lin et al. 2017), (h) FCOS (Tian et al. 2019), (i) ATSS (Zhang et al. 2020b), (j) TOOD (Feng et al. 2021), (k) PAA (Kim and Lee 2020), (l) Ours-Standard, (m) Ours-AMSP-VConv, (n) Ours-AMSP-VConv + NMS Similar.
@@
We adopt AP and AP50 as the primary metrics for model accuracy evaluation, with precision (P) and recall (R) as supplementary indicators. To showcase the generalizability of our network, we trained it on the URPC (Zhanjiang) (Liu et al. 2021) dataset, from the 2020 National Underwater Robotics Professional Competition and the extensive RUOD dataset. The URPC dataset contains 5,543 training images across five categories, with 1,200 images from its B-list answers serving as the test set. The RUOD dataset (Fu et al. 2023) contains various underwater scenarios and consists of 10 categories. It includes 9,800 training images and 4,200 test images.
@@
Table 3: Comparison with existing methods on the URPC and RUOD datasets. Ho: holothurian’s AP50, Ec: echinus’s AP50, St: starfish’s AP50, Sc: scallop’s AP50, Wa: waterweeds’s AP50. AP: AP@[0.5:0.05:0.95], AP50: AP@0.5. Bolding and underlining is highest, underlining only is second-highest.
@@
Ablation of AMSP-VConv:
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (No.62301105), the 2022 National Undergraduate Innovation and Entrepreneurship Training Program Project (No.202210577003), National Key Research and Development Program of China (No.2018AAA0100400), China Postdoctoral Science Foundation (No.2021M701780), the High Performance Computing Center of Dalian Maritime University, and the Supercomputing Center of Nankai University. We are also sponsored by CAAI-Huawei MindSpore Open Fund.